What is the wavelength of the sound emitted by an airplane flying at an elevation of 40,000 feet where the temperature is -52°C? The frequency of the noise from the engines is 99 Hz. Express answer to two decimal places.
Answer #1

the velocity of sound at temperature "T"

V = 331 + 0.6 T

= 331 + 0.6 x ( -52°C )

= 299.8 m/sec

= C/

= C /

= 299.8 / 99

= 3.02828 m

the wavelength of the sound emitted by an airplane is = 3.02828 m